OpenStudy (anonymous):

The high school band has 12 more girls than boys. The band has a total of 106 students. How many boys are in the band?

OpenStudy (anonymous):

2x + 12 = 106 2x = 94 x = 48

OpenStudy (anonymous):

my bad 47 lol

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Let x be the number of girls in the band, then the number of boys is x - 12. As the band has a total of 106 students, we have: x + (x - 12) = 106 Solving the equation for x, we have x + x - 12 = 106 2x = 118 x = 59 So there are 59 girls in the band. 106-12=94 1/2 of 94 is 47 There 47 boys and 59 girls.